# Removing Reflections from RAW Photos

### v2

@@ Abstract
We describe a system to remove real-world reflections from images for consumer photography. Our system operates on linear (RAW) photos, with the (optional) addition of a contextual photo looking in the opposite direction, e.g., using the “selfie” camera on a mobile device, which helps disambiguate what should be considered the reflection. The system is trained using synthetic mixtures of real-world RAW images, which are combined using a reflection simulation that is photometrically and geometrically accurate. Our system consists of a base model that accepts the captured photo and optional contextual photo as input, and runs at 256p, followed by an up-sampling model that transforms output 256p images to full resolution. The system can produce images for review at 1K in 4.5 to 6.5 seconds on a MacBook or iPhone 14 Pro. We test on RAW photos that were captured in the field and embody typical consumer photographs.
Figure 1: Results of our reflection removal system. We use linear (RAW) images with an optional contextual photo, and output full resolution linear clean and reflection images for editing (shown at $2\mathrm{K}$). Prior works use tone-mapped images at about $256^{2}$ pixels, yielding lower quality and less accurate color. The brightness/contrast differences relative to captured photos arise from reflection removal, and are correct for the scenes.
@@ 1 Introduction
Taking pictures through glass is difficult. Light reflects off of the glass and linearly mixes with the subject, creating a distraction. Photos from cars and airplanes show the cabin, photos from buildings include the ceiling lights, photos of paintings are covered by haze, and shots while window shopping are photo-bombed by the photographer themself—to name just a few cases.
Removing undesirable reflections is difficult due to the wide diversity of locations and circumstances in which they occur. Locations include shopping destinations (photos into stores and displays), traveling (photos from cars, trains, airplanes, ferries, and ships), buildings (photos from restaurants, hotels, conference centers, and houses), museums of art, history, and science (photos of paintings, sculptures, and artifacts; aquariums, zoos, and technology), and in special situations (eyeglasses, specular objects, and monitors or screens). In any location, many circumstances influence the reflection: time of day (dusk, dawn, midday, night), lighting (sunlight, overcast, incandescent, tungsten, LED, colored), scene semantics (trees, clouds, scenery; streets, cars, people), illuminant power (bright, dark), and scene appearance (complex textures or simple colors and shapes). These diverse factors introduce priors on reflections because glass is usually placed judiciously in the real world.
One reflection removal technique involves taking a second photo. A black material is placed behind the glass to allow only reflected light to reach the camera. If this reflection image, and the original mixture image, are stored in a format that preserves the linear relationship between pixel values and the scene luminance (e.g., RAW), then these two scene-referred images can be subtracted to recover the image that transmitted through the glass. This transmission image can be recovered because light mixes by addition in photosites on the sensor.
Removal-by-subtraction has been used extensively to create datasets [27, 48], but it fails if there is motion in the reflected scene, including lighting changes, which restricts where data is collected. Alternatively, a glass pane can be manually placed, but the scene and lighting are typically similar on both sides of the glass. As recently noted by Lei [27], training and evaluations on datasets that do not represent real use cases can be misleading.
This paper presents a reflection removal system for consumer photography that addresses the following requirements:
@@
Few previous works satisfy all of these, which affect design and evaluation. In particular, data should represent the expected use for the system. To address req. 1, we accurately synthesize images that obviate the capture of ground truth images that prior systems use. By training on such data, we avoid using datasets that are skewed to unrealistic situations that allow ground truth capture.
Specifically we use linear scene-referred images to simulate images with photometric and colorimetric calibration (e.g., RAW). Thus, if we synthetically combine a (presumed transmitted) image of a storefront with a (presumed reflected) image of sunlit buildings, the reflection will be brighter, typically bluer in white balance, but attenuated by the physically plausible reflectance of glass.
To address reqs. 24, we avoid using video, frame bursts, and stereo pairs. Instead, we accept a single optional contextual photo to help identify what created the reflection. This second photo does not need to be captured simultaneously, or registered with the original photograph. In fact, it could be captured by turning around, pointing away from the window, and snapping the shutter again.
To address req. 5, we design a novel upsampler with a flexible output resolution; upsampling is imperative and non-trivial, but largely disregarded in the literature. To meet req. 6 we output both reflection and transmission images so users can blend them to contend with the inevitable long tail of practical failures.
Contributions: this work describes methods to
@@
    synthesize training images from which models can directly generalize;
@@
    use a contextual photo to identify the reflection;
@@
    remove reflections in about 5 seconds at $1\mathrm{K}$ resolution for review;
@@
    remove reflections at typical full image resolutions.
The paper is organized into prior work (Sec. 2), reflection synthesis (Sec. 3), removal (Sec. 4), and results (Sec. 5); geometric simulation (Appendix BC), data collection (Appendix D), modeling (Appendix E), and results (Appendix F).
@@ 2 Prior work
Reflection synthesis. Prior methods are trained with heuristically mixed pairs of tone-mapped images [9, 20, 16, 31, 55, 47, 56, 21, 6, 11]. Such mixing is inaccurate, so non-linear methods have been used [51, 24]. Physically based methods nonetheless use tone-mapped images [24]. Successful methods however require ground truth images to train models that generalize [27], typically at approximately a a10101010:$1$ ratio of synthetic and real [20, 27, 47, 58, 50, 29, 36, 32]. This ratio raises issues of dataset scale and diversity because ground truth capture is tedious and restrictive. The largest dataset of real images to-date [59] has 14,952 pairs ($10^{4}$), but methods such as [20, 47, 58] require pre-training on much larger datasets which exceed $10^{6}$ (e.g., ImageNet [41]). We synthesize photometrically accurate images so ground truth training images are not needed, and train models from scratch on millions of reflection examples, which improves performance.
Removing high resolution reflections. Most methods operate at about $256^{2}$ pixels, and cannot be trivially scaled up. To be useful, systems must provide preview images at $\approx 1\mathrm{K}$ pixels, and final outputs of $4\mathrm{K}$ and larger. Prasad [36] use a base model at $256^{2}$ pixels, and an upsampler that outputs beyond $4\mathrm{K}$ pixels. Their system is fast, but we show that it re-introduces sharp residual reflections. We design an upsampler with similar performance that removes sharp reflections.
Inference on RAW images. Most prior methods apply reflection removal to $8$-bit display-referred images, such as internet JPEGs. These images have typically been white-balanced, tone-mapped, denoised, sharpened, and compressed. We reframe dereflection as operating on scene-referred (RAW) images. Lei [26] subtract pairs of RAW images to suppress the reflection before converting to $8$-bit for full removal. We operate on RAW end-to-end. RAW inputs improve prior methods, but our system outperforms them.
@@ 3 Reflection synthesis
Our pipeline for removing reflections uses a base model and an upsampler (Sec. 4) that are trained solely on simulated images, which overcomes the scaling bottleneck of needing to capture real reflections. We simulate reflections photometrically by summing pairs of scene-referred images, which are linear with respect to scene luminance. In contrast, images in most $8$-bit formats are display-referred—non-linearly related to luminance. Scene-referred images originate from sensor data stored in RAW format, such as Adobe Digital Negative (DNG). The transformation of RAW data into display-referred images is described by Adobe Camera RAW (ACR), the DNG spec. [1] pp.99-104, and the DNG SDK [1] as follows:
@@
    Linearize (e.g. remove vignetting and spatially-varying black levels)
@@
    Subtract the black level
@@
    White balance [^1] ACR defines two possible paths to a white balanced image—we discuss this below.
@@
Step 8 yields a finished image that can be stored in $8$ bits, but its pixel values are non-linearly related to scene luminance because Step 7 performs an array of proprietary, non-linear, and spatially varying effects that are not usefully modeled with a gamma curve as is often done [58, 52, 29]. Most importantly, realistic reflections cannot be simulated by summing pairs of finished images.
Which earlier step is most appropriate for simulation? The outputs of Steps 5 and 6 are linear, but the illuminant color has been removed by white balancing—accurate reflections cannot be simulated here because scenes that reflect from and transmit through glass are often illuminated by light sources with differing colors, and those colors mix before white balancing. The output of Step 3 is linear, preserves the illuminant color, and has been demosaicked, but colors are with respect to a sensor-specific spectral basis—images from different sensors cannot be summed here. The output of Step 4 is however ideal: the XYZ color space is sensor-independent, yet the illuminant color has not been removed (unlike prior works [2]), and pixels are linear with respect to luminance. We therefore select Step 4 and XYZ color space to simulate photometrically accurate reflections.
Note that the ACR steps (above) differ from many cameras and the literature [2, 7, 22], wherein white balance is applied before converting to XYZ with the forward matrix. ACR also supports that ordering (DNG Spec.[1] p103, matrix FM), but reflection simulation requires the opposite. Fortunately, ACR specifies a second path that uses color matrices (DNG Spec. [1] pp101-103, matrix CM), to transform to XYZ before white balancing. All DNGs are required to provide such color matrices, whereas the forward matrices of the first path are optional. [^2] ACR recommends forward matrices under extreme lighting (DNG Spec. [1] pp.101-103), for which they are more precise. Reflection simulation however requires using color matrices. Both depend on the as-shot illuminant; see Funcs. A10, A5, A8. We therefore use the color matrix path. In Sec. 5, we show that this color processing yields synthetic training data with sufficient realism for models to generalize to photos in-the-wild from other cameras, while prior methods do not. Appendix G and Func. A4 detail how to extract these XYZ images using the DNG SDK [1].
@@ 3.1 Photometric reflection synthesis
Our most fundamental simulation principle is the additive property of light: glass superimposes the light field from a reflection and transmission scene to form a mixture. The resulting mixture image $m=t+r$ accumulates (with equal weight) photons from the transmission scene into a transmission image $t$ and a reflection image $r$. We simulate $t$ and $r$ from images in linear XYZ color (ACR Step 4).
The first photometric property is illuminant color, which often differs between $t$ and $r$ because the glass in consumer photographs typically separates indoor and outdoor spaces. Otherwise, the photographer could walk around the glass to take their photo. Even in specialized scenes like museum display cases, the case is often internally illuminated, making its illuminant color different than in the gallery at large. By representing $(t,r)$ in XYZ color before white balancing, the illuminant colors are mixed.
The second property is the power of light. In typical scenes, the illuminant power differs on either side of the glass ($t$ and $r$ differ in brightness). The number of photons that strike the sensor is scaled by the exposure $e=s\cdot g/n^{2}$, for shutter speed $s$, aperture $n$, and gain $g$ (ISO). We normalize the exposures of $t$ and $r$ so pixels are proportional to scene luminance up to a shared constant. This non-exposed mixture $m^{\prime}$ is $m^{\prime}=t^{\prime}+r^{\prime}$, $t^{\prime}=t/e_{t}$, and $r^{\prime}=r/e_{r}$, for exposures $e_{t}$ and $e_{r}$. We simulate a capture function $\mathcal{C}$ that re-exposes and re-white balances $m^{\prime}$ by exposing the mean pixel to a target value $\tau$, $m=\mathcal{C}(m^{\prime})$, $\mathcal{C}(m^{\prime})=We^{\prime}m^{\prime}$, and $e^{\prime}=\tau/\operatorname{\mathbb{E}}[m^{\prime}]$, where $W$ is a $3\times 3$ matrix that white balances in XYZ (Func. A3, Sec. A.1). If pixels in $t$ or $r$ are saturated, $e^{\prime}=1/\min(\max(t^{\prime}),~{}\max(r^{\prime}))$, to ensure they remain so. Lastly, $m$ is converted to scene-referred, linear RGB to train dereflection models. See Sec. 4, Func. 1 and Appendix A for details.
Mixtures $m$, above, are photometrically mixed, but they are not always useful. When saturation dictates the re-exposure $e^{\prime}$, additional pixels can be clipped, modeling over-exposed $m$. Images $t^{\prime}$ or $r^{\prime}$ can also be so dark that they are invisible, or so mutually destructive that one would struggle to identify the subject. These photos do not model $m$ that photographers care about. We therefore collect a large dataset of images and search for $(t,r)$ that yield well exposed and mixed $m$. This search introduces photometric and semantic priors on $m$, $t$, and $r$ (e.g., skies often create reflections). See Appendix D for details.
@@ 3.2 Geometric reflection synthesis
Our second fundamental simulation principle is that we want the mixtures to be geometrically valid. Specifically, denoting the synthetic images to be added together as $(t,r)$ and our source image pairs as $(i,j)\in\mathcal{D}$, we synthesize $t=T(i)$ and $r=R(j)$ by modeling Fresnel attenuation, perspective projection, double reflection, and defocus. We omit from $T$ effects related to global color, dirt, and scratches, since photo editing tools can correct them. We also model a physically calibrated amount of defocus blur, and find that most reflections are sharp as also noted in [27]. See Appendix B for details.
@@ 3.3 The contextual photo
We accept an optional contextual photo $c$ that directly captures the reflection scene to help identify the reflection $r$. Capture of $c$ can be simultaneous with the secondary front camera (selfie) on a mobile device, or briefly later. We make three observations about the views of $c$ and $r$ (see Fig. A11):
@@
    If the front camera is used, the reflection scene might be partially occluded by the photographer.
Image $c$ will therefore often contain little matching content unless it is captured carefully. We minimize such burden by allowing $c$ to be any view of the reflection scene. Crucially, this relaxation also facilitates geometric simulation. We scalably model $c$ by cropping source images into a disjoint left/right half (or top/bottom). The contextual image encodes information about the lighting and scene semantics because we use the same capture function $\mathcal{C}$ with the same white balance as $(m,t,r)$. See Sec. 3.1, Func. 1, and Appendix C for details.
@@
2:  Split $i$ similarly: randomly select a transmission part $t$ and discard the other.
@@
6:  Compute a new exposure $e$ for $m$. {Func. A2}
7:  Compute a white balance transform XYZ_to_XYZ_awb from $e\cdot m$. {Func. A3}
8:  White balance $m$ by applying XYZ_to_XYZ_awb.
@@
10:  Get the transform XYZ_D50_to_sRGB. {SDK Func. A13}
11:  Transform $(m,t,r,c)$ to linear sRGB using XYZ_D50_to_sRGB.
@@
Figure 2: The base model. Mixture and context images are projected into a high dimensional space using a shared backbone [44] (labeled P-Net; weights are shared), and a feature fusion network [45] (labeled F-Net; weights are shared). The context features are used to predict affine transforms for each feature channel at each resolution. Channel-wise modulation is used because contextual photos do not always include content that can be matched. Modulation can help identify the reflection in the feature space. We use two conv-mod-deconv operations of [23] within the modulated merge blocks. The FPN combine op is a fast normalized fusion module from the BiFPN architecture [45].

### v3

@@ Abstract
We describe a system to remove real-world reflections from images for consumer photography. Our system operates on linear (RAW) photos, and accepts an optional contextual photo looking in the opposite direction (e.g. the “selfie” camera on a mobile device). This optional photo disambiguates what should be considered the reflection. The system is trained solely on synthetic mixtures of real RAW photos, which we combine using a reflection simulation that is photometrically and geometrically accurate. Our system comprises a base model that accepts the captured photo and optional context photo as input, and runs at 256p, followed by an up-sampling model that transforms 256p images to full resolution. The system produces preview images at 1K in 4.5-6.5s on a MacBook or iPhone 14 Pro. We show SOTA results on RAW photos that were captured in the field to embody typical consumer photos, and show that training on RAW simulation data improves performance more than the architectural variations among prior works.
Figure 1: Results of our reflection removal system. We use linear (RAW) images with an optional contextual photo, and output the clean and reflection images in linear color for editing, at full resolution (shown at $2\mathrm{K}$). Prior works use tone-mapped images at $\approx 256$p, yielding lower quality and inaccurate color. Brightness/contrast changes relative to captured photos arise from reflection removal, and are correct.
@@ 1 Introduction
Taking pictures through glass is difficult. Light reflects off of the glass and linearly mixes with the subject, creating a distraction. Photos from cars and airplanes show the cabin, photos from buildings include the ceiling lights, paintings are covered by haze, and window shopping shots are photo-bombed by the photographer—to name just a few cases.
Removing unwanted reflections is difficult because they occur in a diverse range of locations and situations. Locations include shopping spots, traveling (cars, planes), buildings, museums, and special cases (eyeglasses, screens). Reflections depend on the time, lighting (e.g. incandescent), scene (trees, streets), illuminant power, and appearance (complex textures or simple shapes). These factors create priors because glass is placed carefully in the world.
One way to remove reflections involves capturing a second photo with a black material placed behind the glass to allow only reflected light to reach the camera. If this reflection image, and the original mixture image, are stored in a format that preserves the linear relationship between pixel values and the scene luminance (e.g., RAW), then these two scene-referred images can be subtracted to obtain the image that transmitted through the glass. This transmission image can be recovered because light mixes by addition in photosites on the sensor. Subtraction has been used for datasets [28, 48], but fails under motion or lighting changes; this severely restricts data collection. Alternatively one can place glass panes in the scene, but the scene and lighting are typically similar on both sides. Training and evaluating on such unrealistic data can be unhelpful and misleading.
This paper presents a reflection removal system for consumer photography that targets the following requirements:
@@
Few prior works satisfy these, which affect design and evaluation. In particular, data should match how the system will be used. For req. 1, one needs a large dataset of realistic photos. Prior works require ground truth photos, but capturing them restricts the dataset size and diversity. We synthesize realistic and diverse photos in large quantities.
We synthesize reflections by combining, for example, an image looking at a storefront with one of sunlit buildings presumed to be behind the photographer. To make the synthesis accurate, we use linear scene-referred images with known photometric and colorimetric calibration, and combine them in a physically correct way. For example, the reflected buildings are typically brighter and bluer than the storefront, but will be attenuated by reflection off the glass.
To address req. 2 (capture time) and 4 (processing time), we avoid asking the user to capture video, bursts of frames, or stereo photos. These help identify what created the reflection, but they slow down processing. Instead, we allow the user to capture an optional, contextual photo. This photo does not need to be captured simultaneously, or registered with the original photo. In fact, it could be captured by quickly turning around and looking away from the window.
To address req. 5, we use a novel upsampler with a flexible output resolution. Note that upsampling is imperative and non-trivial, but mostly disregarded in the literature. To meet req. 6 we output reflection and transmission images so users can remix them to fix the long tail of practical failures.
Contributions. In this work, we
@@
    show how to synthesize training data such that models do not need to be fine-tuned on real images;
@@
    show that training/testing on RAW improves performance significantly—more than prior model variations;
@@
    use a contextual photo to help identify the reflection;
@@
    significantly reduce upsampling artifacts while producing output at 1K in 5s for review, and at full resolution.
This paper is best read with hyperlinks into the supplement. See arXiv or the project website for a complete version. Prior work is outlined in Sec. 2, reflection synthesis (Sec. 3), removal (Sec. 4), and results (Sec. 5). In supplemental sections we discuss simulation (Appendix AC), data collection (Appendix D), modeling (Appendix E), and results (Appendix F).
@@ 2 Prior work
Reflection synthesis. Prior methods are trained with heuristically mixed pairs of tone-mapped images [9, 20, 16, 31, 56, 47, 57, 21, 6, 11]. Such mixing is inaccurate, so non-linear methods have been used [52, 24]. Physically based methods nonetheless use tone-mapped images [24]. Successful methods however require ground truth images to train models that generalize [28], typically at approximately a $10$:$1$ ratio of synthetic and real [20, 28, 47, 59, 51, 29, 35, 32]. This ratio raises issues of dataset scale and diversity because ground truth capture is tedious and restrictive. The largest dataset of real images to-date [60] has 14,952 pairs ($10^{4}$), but methods like [20, 47, 59] require pre-training on datasets larger than $10^{6}$ (e.g., ImageNet [40]). We synthesize photometrically accurate images to obviate ground truth training images, and train models from scratch on more than 1M examples, which improves performance.
Removing high resolution reflections. Most methods operate at $\approx 256^{2}$ pixels, and cannot be trivially scaled up. Useful systems must create preview images at $\approx 1\mathrm{K}$ pixels, and final outputs beyond $4\mathrm{K}$. Prasad [35] use a base model at $256^{2}$ pixels, and an upsampler that yields $\geq 4\mathrm{K}$ pixels. Their fast upsampler re-introduces sharp reflections. Our upsampler is similarly fast, but removes sharp reflections.
Inference on RAW images. Most prior methods apply reflection removal to $8$-bit display-referred images, such as internet JPEGs. Such images have been white-balanced, tone-mapped, denoised, sharpened, and compressed. We reframe dereflection to operate on scene-referred (RAW) images. Lei [26] subtract pairs of RAW images to suppress the reflection before converting to $8$-bit for full removal. We operate on RAW end-to-end. RAW inputs improve prior methods, but our system outperforms them.
@@ 3 Reflection synthesis
Our pipeline for removing reflections uses a base model and an upsampler that are trained solely on simulated images (Sec. 4), which overcomes the scaling bottleneck of needing to capture real reflections. We simulate reflections photometrically by summing pairs of scene-referred images, which are linear with respect to scene luminance. In contrast, images in most $8$-bit formats are display-referred—non-linearly related to luminance. Scene-referred images originate from sensor data stored in RAW format, such as Adobe Digital Negative (DNG). The transformation of RAW data into display-referred images is described by Adobe Camera RAW (ACR), the DNG spec. [1] pp.99-104, and the DNG SDK [1] as follows:
@@
    Linearize (e.g. remove vignetting and black levels)
@@
    Subtract the scalar black level
@@
    White balance [^1] ACR defines two white balancing paths, and we leverage one that differs from many cameras and the literature [3, 7, 22]. In the literature, white balance is applied before converting to XYZ with the forward matrix. ACR also supports that ordering (DNG Spec.[1] p103, matrix FM), but reflection simulation requires the opposite (as explained in Sec. 3). Fortunately, ACR specifies a second path that uses color matrices (DNG Spec. [1] pp101-103, matrix CM), to transform to XYZ before white balancing. All DNGs are required to provide such color matrices, whereas the forward matrices of the first path are optional. ACR recommends forward matrices under extreme lighting (DNG Spec. [1] pp.101-103), for which they are more precise. Both paths however depend on the as-shot illuminant; see ACR Funcs. G, G, G. In Sec. 5, we show that this color processing yields synthetic training data with sufficient realism for models to generalize to photos in-the-wild from other cameras, while prior methods do not.
@@
Step 8 yields an $8$-bit finished image for storage, but its pixel values are non-linearly related to scene luminance because Step 7 performs proprietary, non-linear, and spatially varying effects that cannot be modeled with a gamma curve as is often done [59, 53, 29]. Realistic reflections therefore cannot be simulated by summing pairs of finished images.
Which earlier step is most appropriate for simulation? The outputs of Steps 5 and 6 are linear, but the illuminant color has been removed by white balancing—accurate reflections cannot be simulated here because scenes that reflect from and transmit through glass are often illuminated by light sources with differing colors, and those colors mix before white balancing. The output of Step 3 is linear, preserves the illuminant color, and has been demosaicked, but its colors are with respect to a sensor-specific spectral basis—images from different sensors cannot be summed here. The output of Step 4 is however ideal: the XYZ color space is sensor-independent, the illuminant color is preserved (unlike prior works [3]), and pixels are linear with respect to luminance. We therefore select Step 4 and XYZ color space to simulate photometrically accurate reflections.
@@ 3.1 Photometric reflection synthesis
Our most fundamental simulation principle is the additive property of light: glass superimposes the light fields from a reflection and transmission scene to form a mixture. The resulting mixture image $m=t+r$ accumulates (with equal weight) photons from the two scenes into a transmission image $t$ and a reflection image $r$. We simulate $t$ and $r$ from images in linear XYZ color (ACR Step 4).
The first photometric property is the illuminant color, which often differs between $t$ and $r$ because the glass in consumer photographs typically separates indoor and outdoor spaces. Otherwise, the photographer could walk around the glass to take their photo. Even in specialized scenes like museum display cases, the case is often internally illuminated, making its illuminant color different than in the gallery at large. By representing $t$ and $r$ in XYZ color before white balancing, the illuminant colors are mixed.
The second property is the illuminant power. In typical scenes, this power differs on either side of the glass ($t$ and $r$ differ in brightness). The number of captured photons is scaled by the exposure $e=s\cdot g/n^{2}$, for shutter speed $s$, aperture $n$, and gain $g$ (ISO). We normalize the exposures of $t$ and $r$ by $e$ so pixels are proportional to scene luminance up to a shared constant. This un-exposed mixture is $m^{\prime}=t^{\prime}+r^{\prime}$, $t^{\prime}=t/e_{t}$, and $r^{\prime}=r/e_{r}$, for exposures $e_{t}$ and $e_{r}$. We simulate a capture function $\mathcal{C}$ that re-exposes and re-white balances $m^{\prime}$ by exposing the mean pixel to a target value $\tau$, $m=\mathcal{C}(m^{\prime})=We^{\prime}m^{\prime}$, and $e^{\prime}=\tau/\operatorname{\mathbb{E}}[m^{\prime}]$, where $W$ is a $3\times 3$ matrix that white balances in XYZ (Func. A.1, Sec. A.1). If pixels in $t$ or $r$ are saturated, $e^{\prime}=1/\min(\max(t^{\prime}),~{}\max(r^{\prime}))$, to ensure they remain so. Lastly, $m$ is converted to scene-referred, linear RGB to train models.
The full simulation is described in Func. 3.3 and Appendix A. This function produces mixtures $m$ are photometrically accurate, but they aren’t always useful. When saturation dictates the re-exposure $e^{\prime}$, pixels can be clipped, modeling over-exposed $m$. Images $t^{\prime}$ or $r^{\prime}$ can also be so dark that they are invisible, or so mutually destructive that one would struggle to identify the subject. These photos do not model $m$ that photographers care about. We therefore collect a large dataset of images and search for well exposed and well mixed $m$. This search introduces photometric and semantic priors on $m$, $t$, and $r$ (e.g., skies often reflect). See Appendix D.
@@ 3.2 Geometric reflection synthesis
Our second fundamental simulation principle is that mixtures must be geometrically valid. Denoting the images to be summed as $t$ and $r$, and our source image pairs as $(i,j)$, we synthesize $t=T(i)$ and $r=R(j)$ by modeling spatially varying Fresnel attenuation, perspective, double reflection, and defocus. We omit from $T$ the effects of global color, dirt, and scratches; editing tools can correct them. We model a physically calibrated amount of defocus blur; most reflections are sharp as also noted in [28]. See Appendix B.
@@ 3.3 The contextual photo
We accept an optional contextual photo $c$ that directly captures the reflection scene to help identify the reflection $r$. Capture of $c$ can be simultaneous with the secondary front camera (selfie) on a mobile device, or briefly later. We make three observations about the views of $c$ and $r$ (see Fig. S3):
@@
    If the selfie camera is used, the reflection scene might be partially occluded by the photographer.
Image $c$ will therefore often contain little content that matches with $r$ unless it is captured carefully. We avoid placing such a large burden on the user, and allow them to capture any view, $c$, of the reflection scene. Crucially, this relaxation also facilitates the geometric simulation. We scalably model $c$ by cropping source images into a disjoint left/right half (or top/bottom). The context image encodes information about the lighting and scene because we use a capture function $\mathcal{C}$ with the same white balance as $(m,t,r)$. See Sec. 3.1, Func. 3.3, and Appendix C for details.
@@
2:  Split $i$ similarly: randomly select a transmission part $t$.
@@
6:  Compute a new exposure $e$ for $m$. {Func. A.1}
7:  Compute WB matrix XYZ_to_XYZ_awb. {Func. A.1}
8:  White balance (WB) $m$ by applying XYZ_to_XYZ_awb.
@@
10:  Get the transform XYZ_D50_to_sRGB. {SDK Func. G}
11:  Transform $(m,t,r,c)$ to linear sRGB.